A spherical tank has a radius of 8 m. If the tank is completely filled with water, compute the work required to pump all of the water to the level of the top of the tank.
Answer #1

8 Centre q raduis of -Y Equation of the circle assuming the the tank to lie at origin x² + y²=82 x = 164 - yz ie equal to cro8 W = 980017 512 - 842-644 + y3) dy -- 8 co W 98001T 512y 8y! - 649? + y 4 -8 W=9800T [(-şlar zularin 8)3- 32(8) 84 $12 (-0)